OpenStudy (supercait10):

1. If Q=2, Evaluate (34 + 18. q) - (5q + 7) A. 38 B. 79 C. 31 D. 21

OpenStudy (anonymous):

\[(34+18\times 2)-(5\times 2+7)\]is a better start

OpenStudy (anonymous):

then maybe a calclulator

OpenStudy (anonymous):

actually get none of the above could there be a typo there?

OpenStudy (supercait10):

nope that's the choices

OpenStudy (anonymous):

can you post a screenshot of the question?

OpenStudy (anonymous):

maybe i read "18.q" wrong not really clear there

OpenStudy (supercait10):

yea i'll get a screenshot hold on

OpenStudy (anonymous):

ok

OpenStudy (supercait10):

the last one

OpenStudy (triciaal):

why is this same question posted again?

OpenStudy (supercait10):

Because my teacher gave me a second chance at this question and he said if I couldn't figure it out after five tries he said I can ask for help so I am

OpenStudy (anonymous):

Ah 5 to the power of q, that means the following \[5^2 = (5)(5) =25\]

OpenStudy (anonymous):

oooh !!

OpenStudy (anonymous):

nothing like the equation editor

OpenStudy (triciaal):

but we can't help you if the question is not clear.

OpenStudy (anonymous):

Please use "^" to indicate to the power of next time :)

OpenStudy (supercait10):

i just posted a file showing the test look at it to understand it

OpenStudy (anonymous):

\[(34+18\times 2)-(5^ 2+7)\]

OpenStudy (anonymous):

i guess you are supposed to be practising the "order 'o operations" right?

OpenStudy (anonymous):

you know what to do first?

OpenStudy (supercait10):

yup now starting Algebra and haha no my weakest link in school is math I barley passed Math last year so this is beyond me I better understand History then i do math

OpenStudy (anonymous):

square first

OpenStudy (anonymous):

i.e. compute \(5^2=5\times 5\) don't think too hard here

OpenStudy (anonymous):

what do you get?

OpenStudy (supercait10):

hold on had to get a drink

OpenStudy (anonymous):

yeah i need one too maybe some bulleit rye

OpenStudy (supercait10):

lol maybe

OpenStudy (anonymous):

did you manage to compute \(5\times 5\) yet?

OpenStudy (supercait10):

yea

OpenStudy (anonymous):

care to share it with the class?

OpenStudy (supercait10):

yea

OpenStudy (anonymous):

...

OpenStudy (supercait10):

25

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

so now \[(34+18\times 2)-(5^ 2+7)=(34+18\times 2)-(25+7)\]

OpenStudy (anonymous):

now lets do the multiplication \(18\times 2=?\)

OpenStudy (supercait10):

36

OpenStudy (anonymous):

ok so now we are at \[(34+36)-(25+7)\] add in each parentheses what do you get?

OpenStudy (supercait10):

(34 + 36 = 70) (25 + 7 = 32) all together = 102

OpenStudy (anonymous):

there is no such mathematical operation as "all together" is there?

OpenStudy (anonymous):

you should be looking at \[70-32\]

OpenStudy (supercait10):

i'm tired so I'm really not paying attention to what I type

OpenStudy (anonymous):

then go to bed and do this arithmetic some other day

OpenStudy (supercait10):

i can't I have to have this submitted by 11:30 tonight or I fail it